A flight attendant wants to change the temperature of the air in the cabin from 18.0°C to 23.9°C without changing the pressure. What fractional change in the number of moles of air in the cabin would be required?
Answer #1

AV=nRT T = 18°C - 273+ 18 = 2.9k T2 = 23.9°C = 273+23.9 T₂ = 296.9k PV = n RT NT, = constant na ne-ni - 1 - - 1 = T Ti ni o T